give nitroprusside 5mcg/kg/minute via continuous infusion for a patient weighting 205 lbs. nitroprusside is available in a solution of 200mg in 250mg dsw. what rate on the infusion pump will deliver the correct dose
Answer #1

ANSWER.

* Calculation of rate of infusion.

         Rate(ml/hr) = Volume(ml) /Time(hr)

      Formula for calculation of volume = Desired dose/Dose in hand * Quantity.

    Here

         Ordered dose = 5 mcg/kg/min.

        Patient weight = 205 lbs.

    Convert lbs to kg for calculation

            1 lbs = 0.454 kg.

        Patient weight (kg) = 205 * 0.454

                                       = 93.07 kg.

         Desired dose (D) = 5 * 93.07

                                      = 465.4 mcg

       Convert mcg to mg for calculation

                   1 mcg = 0.001 mg.

          Desired dose(D) = 465.4 * 0.001

                                      = 0.47 mg/min.

          Dose in hand (H) = 200 mg.

                 Quantity (Q) = 250 ml.

    As per above formula

                   Volume of medication needed = 0.47 /200 * 250

                                                                    = 0.002 * 250

                                                                    = 0.5 ml/min.

                  Convert ml/min to ml/hr by multiply with 60.

                              Rate (ml/hr) = 0.5 ml/min * 60

                                                   = 30 ml/hr.